OpenStudy (anonymous):

Find the lowest common denominator that would eliminate the fractions in this equation. x/2+1/4=3/5

Parth (parthkohli):

20 is the answer.

OpenStudy (anonymous):

x/2=3/5-1/4 x/2=7/20 x=7/10.

OpenStudy (anonymous):

I thought that 20 wasnt the right answer but whats the reduced

OpenStudy (anonymous):

so you have to elimate the fraction so multiply with 20 to remove fraction ans will be 20.

OpenStudy (anonymous):

10 is the right answer